Weekly meeting on MDC readout upgrade

-- BurkhardKolb - 08 Sep 2009


BurkhardKolb, JanMichel, JoernWuestenfeld, KathrinGoebel, MichaelTraxler, SergeyYurevich

Kathrins report

  • see pdf with measurements 1 to 8 boards.
  • one noisy board (b) was found
  • the other boards were all below 1 permille when adjusting the thresholds to about 50 (no individual tuning)
  • looks good
  • still needed is the CTS measurement
  • report to Christian and decision!
  • one new CTS cable at plane 3 sector 4 right side
  • 1 MHz with function generator - some noise data were collected with varying threshold
  • check with analog out for crosstalk/induced noise
  • many CTS cables are produced (except for sector 1)
  • measurements with simulated CTS were made see here

Jans report

  • ethernet is not working on the hub (long packets)
  • all OEP Voltages were measured by the internal ADCs
  • look good with some exceptions were tuning/calibration of the ADC is required ( voltage divider).

Current measurements

  • see first measurements here: MdcAddonCooperation
  • proposal to make test with pulser firing all (or half of all) channels, then increase rate from 1 to 20 kHz (with reasonable thresholds)
  • check rise of current (should stay below 1 A for long MBO)
  • last OEP design leaves uninitialzed MBO at low current
  • there are doubts that the threshold setting worked, has to be checked with proper init file script.

Voltage distribution board/ Fanpw2

  • first board exists and is tested by Torsten
  • changing return currents ( from 5V) influence ground for 3+-3V on daughterboards because they are unregulated on the OEP
    • solution 1: new OEP with regulation for +-3 V (problem: neg low power LDOs not available)
    • solution 2 (favoured): common ground cable: make test with massive copper ground cable
  • Fanpw3 will be done ( 3 months )

latest tests in the cave

  • with 8 OEPBs and daq was done successfully


  • work has started including init from DB
  • Joern will come to GSI mid July to work with Ilse and Attilio
  • link to DataStructure
  • link to OEPNetworkAddresses
  • Michael will provide perl script for daq-anal, sniff, decode...
  • see first tests
  • done


  • It works - but
  • transfer too slow (FPGA to TRB?)
    • at GSI it runs with 75kHz
    • what's wrong at FZD?
    • no news yet
  • edge triggered data reception in FPGA
    • double and triple hits on same address, 0-hits
    • on all MBOs?
    • only on chains of 2 or more MBOs?
  • GDE signal derived from trigger, but MBO derived from CMS
    • GDE to late?
  • GDE release during calibration
    • reset hit registers to early?
    • data ok?
  • Make tests with pulser to measure electronics efficiency
  • new tests with pulser LVL1 up to 1MHz
  • high threshold with EB up to 20kHz - maybe the network with 10Mbits is the problem...
  • shielding - not clear if efficient, new residual plots needed
    • see FZR results (added by CM)
    • my (CM) interpretation of the recent results: NO significant change of the noise level, i cannot recommend to install these additional shieldings, also due to the unclear influence of the HV current, reported by Jrn.
  • charging up has to be investigated

decision on priority was taken

  • priority
    • 1. switch fast to OEP
    • (fix all errors on current design)

current work at GSI

  • Sector 4, layer 3 (POFs) and 4 started
  • all cables and POFs should be fixed to chamber, not to supports
  • new cable type for Common Stop needed (Michael)
  • better labels needed (Kathrin)

OEP work

  • Attilio reported about bit shifts in the FPGA
    • problem with Lattice design software
    • is manufacturer contacted? yes
    • problem not yet understood
    • comes and goes
    • now has disappeared
  • one link is now working
    • with data and calibration triggers
    • short and long MBO working
    • data file with current data format needed
    • new data format (2 times in 1 32-bit word) not yet
  • Jan is working on a 16 link version
    • ID info of several OEPs read out on hub
    • 8 links seem to work
    • problems with place and route software when more than 8 ports are used.
    • send problematic design to Lattice for help
    • still bit errors in media interface, sometimes dependend on recompilation of design, sometimes intermittent in time.
  • patched TRB needed
  • power distribution board to be tested (Kathrin) (5V, +-3V, 1.4V, 3.5V)
  • power now supplied by Genesys power supplies
  • test in cave with CMS needed for general green light
  • LVDS CMS signal cabled up
  • cross talk can be seen up to a distance of about 5mm
  • watch out during cabling!!!
  • needs now realistic cabling with more MBOs
  • test in cave with 1 MBO and 1 OEPB was done, see pictures of noise TestsinCave
  • qfile_009.tif: noise on analog readout during readout
  • qfile_010.tif: noise on analog readout during readout
  • qfile_011.tif: noise on analog readout during readout
  • readout test was running whole night through
  • agreement was reached on the possibility to switch between two different compression formats for debugging
  • work starts now (SY) to get it in standard readout to HLD-files

New transmission problem 21. July

  • Attilio reworked the FPGA code to support debugging
  • sometimes empty events with some wordcount are sent
  • problem seems to be in sending part
  • has to be solved until next week
  • solved

How to load thresholds and firmware

  • manipulate threshold in Oracle, output to files
  • automatic load via TRBnet to TDCs
  • take threshold data from FPGA after update and insert into data stream?

efficiency measurements at FZD

  • starts at 1900V (plateau), not understood yet, maybe gas mixture?
  • try out Argon isobutan mix
  • change gas flow (oxygen?)
  • small gas leaks are fixed, but no significant improvement at 60/40 mix
  • plateau at 1650 with 80/20 mix, see efficiency plateau
  • new measurements with gas flow switched off indicate leakage and contamination (oxygen) when flow is too low
  • He leak detection will be performed again
  • O2 measurements at FZD, manual available now
  • electrolyte cell was empty
  • needs new evaluation

I Attachment Action Size Date Who Comment
CellvsT1_s2m0l0.jpgjpg CellvsT1_s2m0l0.jpg manage 43 K 2009-03-31 - 12:49 JoernWuestenfeld Cell vs. Time1 for layer0
CellvsT1_s2m0l1.jpgjpg CellvsT1_s2m0l1.jpg manage 37 K 2009-03-31 - 12:50 JoernWuestenfeld Cell vs. time1 for layer1
CellvsT1_s2m0l2.jpgjpg CellvsT1_s2m0l2.jpg manage 41 K 2009-03-31 - 12:50 JoernWuestenfeld Cell vs. time1 for layer 2
CellvsT1_s2m0l3.jpgjpg CellvsT1_s2m0l3.jpg manage 43 K 2009-03-31 - 12:50 JoernWuestenfeld Cell vs. time1 for layer3
CellvsT1_s2m0l4.jpgjpg CellvsT1_s2m0l4.jpg manage 42 K 2009-03-31 - 12:51 JoernWuestenfeld Cell vs. time1 for layer4
CellvsT1_s2m0l5.jpgjpg CellvsT1_s2m0l5.jpg manage 46 K 2009-03-31 - 12:51 JoernWuestenfeld Cell vs. time1 for layer5
Efficency.jpgjpg Efficency.jpg manage 36 K 2009-04-20 - 09:00 JoernWuestenfeld Efficency of variouse gas mixtures and flows
MB-Shielding-FZR.jpgjpg MB-Shielding-FZR.jpg manage 42 K 2009-04-28 - 17:29 ChristianMuentz MB Shielding
MDC-TRB-dataword01.pdfpdf MDC-TRB-dataword01.pdf manage 23 K 2009-04-20 - 14:49 JoernWuestenfeld MDC dataword comparison SAM -> OEP
MDC_DAQ_txtEXT MDC_DAQ_txt manage 1 K 2009-08-04 - 08:45 JoernWuestenfeld Mail from Sergey
MdcGas.jpgjpg MdcGas.jpg manage 95 K 2009-04-21 - 11:40 ChristianMuentz # of cluster He/Isob Ar/Isob
Noise.pdfpdf Noise.pdf manage 15 K 2009-04-22 - 11:06 JoernWuestenfeld Comparing absolute Noise distributions
NoiseComparison.pdfpdf NoiseComparison.pdf manage 105 K 2009-04-06 - 08:45 JoernWuestenfeld Compairing noise situation with and without MBO shielding.
Normalized.jpgjpg Normalized.jpg manage 46 K 2009-04-20 - 09:00 JoernWuestenfeld Efficency normalized to coincidence countrate and fitted with fermi function
PresentationOcc2009-08-25.pdfpdf PresentationOcc2009-08-25.pdf manage 214 K 2009-08-25 - 12:56 KathrinGoebel Occ measurents 2009-08-25
PresentationOcc2009-09-08.pdfpdf PresentationOcc2009-09-08.pdf manage 845 K 2009-09-08 - 12:53 KathrinGoebel Update on Presentation Occ 2009-08-25
efficencyPlateau.jpgjpg efficencyPlateau.jpg manage 31 K 2009-04-07 - 15:00 JoernWuestenfeld Plateau curve of efficency
event_oep.txttxt event_oep.txt manage 709 bytes 2009-08-18 - 11:10 JoernWuestenfeld Event dump with 8 OEP's
gasdep.jpgjpg gasdep.jpg manage 31 K 2009-04-21 - 12:50 JoernWuestenfeld gas dependency plot
noise_213-216313314_thr30.pdfpdf noise_213-216313314_thr30.pdf manage 13 K 2009-08-18 - 08:14 JoernWuestenfeld Noise picture from 8 MBO's
noise_213_thr30.pdfpdf noise_213_thr30.pdf manage 13 K 2009-08-18 - 08:13 JoernWuestenfeld Noise picture from one MBO
noisedifference.jpgjpg noisedifference.jpg manage 25 K 2009-04-22 - 11:06 JoernWuestenfeld Difference in Noise between shielded and unshielded MBO's
overview_occ_140809.pdfpdf overview_occ_140809.pdf manage 6 K 2009-08-18 - 12:29 KathrinGoebel Dataword Occ, 14.08.2009, work in progress
overview_occ_140809_zoom.pdfpdf overview_occ_140809_zoom.pdf manage 5 K 2009-08-18 - 12:30 KathrinGoebel Dataword Occ Detail, 14.08.2009, work in progress
qfile_009.tiftif qfile_009.tif manage 14 K 2009-06-16 - 13:19 BurkhardKolb noise on analog readout during readout
qfile_010.tiftif qfile_010.tif manage 12 K 2009-06-16 - 13:19 BurkhardKolb noise on analog readout during readout
qfile_011.tiftif qfile_011.tif manage 15 K 2009-06-16 - 13:20 BurkhardKolb noise on analog readout during readout
streamer_tubes_nima_1982_Seite_04_a.jpgjpg streamer_tubes_nima_1982_Seite_04_a.jpg manage 61 K 2009-04-21 - 08:17 LotharNaumann Ar/isobuthan dependend working characteristics
voltages_200908211552.txttxt voltages_200908211552.txt manage 2 K 2009-08-25 - 12:56 KathrinGoebel voltages for occ measurements
Topic revision: r45 - 2009-09-08, BurkhardKolb
Copyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki Send feedback | Imprint | Privacy Policy (in German)